OpenStudy (anonymous):

Which is an equivalent expression to negative thirty-three plus positive twenty-four? positive thirty-three minus negative twenty-four negative twenty-four plus positive thirty-three positive twenty-four minus negative thirty-three negative thirty-three minus negative twenty-four

jimthompson5910 (jim_thompson5910):

-33 + 24

jimthompson5910 (jim_thompson5910):

Since the signs are different, think of a subtraction problem (large - small) The result will be negative since the larger number is negative

jimthompson5910 (jim_thompson5910):

For instance -10 + 7 = -3 since 10 - 7 = 3 and this is negative (because the 10 is larger than the 7 and the 10 is negative)

jimthompson5910 (jim_thompson5910):

Oh I see what it's asking now. Here's a hint: subtracting a negative is the same as adding

jimthompson5910 (jim_thompson5910):

Example: 10 - (-7) = 10 + 7

OpenStudy (anonymous):

That Helped! thanks!

jimthompson5910 (jim_thompson5910):

you're welcome